# Leveraging Analog Neutral Atom Quantum Computers for Diversified Pricing in Hybrid Column Generation Frameworks

### v1

@@ I Introduction
Quantum algorithms have been heavily investigated for their potential to deliver speedups for certain classes of combinatorial optimization (CO) problems [1, 2, 3, 4, 5, 6, 7]. These problems involve selecting the optimal solution from a finite, yet exponentially large set of possible solutions subject to constraints. They are ubiquitous in domains such as logistics, scheduling, and resource allocation, where even marginal improvements in optimization can yield substantial operational and economic benefits. Their importance has led to the field of operations research, a domain devoted to the development of applied mathematical techniques to accelerate solvers.
@@
While quantum computing offers alternative algorithmic strategies based on entanglement and superposition, it operates in a large Hilbert space of size $O(2^{N})$, with $N$ the number of qubits in a quantum system. Quantum algorithms must therefore manipulate quantum information in this space, crafting interference patterns so as to concentrate probability on the answer. An example of this recipe for unstructured search is Grover’s algorithm [3], which offers a quadratic speedup over a brute force search. However, such digital algorithms for CO mostly remain out of reach in the noisy intermediate-scale quantum (NISQ) regime due to the prohibitively large number of coherent operations they require.
@@
Quantum annealing approaches, inspired by the adiabatic theorem [8, 9, 10], have been explored extensively as more NISQ-friendly protocols. They typically start by representing an optimization problem as a quadratic unconstrained binary optimization (QUBO), where constraints become slack variables and Lagrange multipliers [11, 12]. Through the annealing, one carries a known product state under a time-varying Hamiltonian such that the final Hamiltonian encodes the cost function (i.e. the QUBO). In the adiabatic limit, the final state will be the solution to the QUBO, although that limit is generally unachievable for NISQ devices with a finite coherence time. Instead, annealing is done in a finite time, and one typically encounters a very small energy gap which leads to the proliferation of defects [13, 14]. Even in a fully coherent system, these defects may lead to observed states that are not even feasible solutions [15]. The quantum approximate optimization algorithm (QAOA) [16, 17], a heuristic method designed to approximate optimal annealing paths on digital quantum computers, suffers from similar problems, although methods such as subspace embedding [18, 19] can remedy them, at the cost of deeper circuits.
Classical algorithms, such as branch-and-price-and-cut [20], harness the structure of the constrained optimization problem for increased performance. They use it to refine heuristics, find cuts, generate new variables, set new bounds. On the other hand, quantum approaches such as those presented above rely mostly on unstructured search. In this dichotomy lies the need for quantum algorithms that take advantage of efficient mathematical formulations, while tackling the biggest bottlenecks with new quantum protocols. Quantum column generation (QCG) schemes [21, 22, 23, 24] fill this need.
@@
They leverage the classical column generation (CG) framework, where the problem is split into an easily solvable master problem and hard subproblems with a common structure. The subproblems focus on identifying new columns (or variables) that may improve the master problem’s optimum. These subproblems are themselves hard CO problems that are solved iteratively many times by heuristics, representing a major bottleneck for performance. Initial results suggest that quantum approaches to the subproblems can lead to better overall results than other heuristics, such as genetic algorithms [22], simulated annealing [23, 21], and even exact methods [23, 21].
@@
In some particular cases, such as the graph coloring problem, the subproblems’ structure as maximum weighted independent sets (MWIS) instances makes them particularly amenable to neutral atom quantum computers (NAQC) [21]. This is because the analog and continuous control of these interacting many-body systems can reach a dynamical regime dominated by the Rydberg blockade, which forbids atoms closer than a certain distance from being both in their respective excited states. This is exactly the independence constraint on edges in a MWIS instance: $n_{i}+n_{j}\leq 1$ ($n_{i}$ is a projector on the Rydberg excited state). Another practically important CO problem, the fleet assignment problem, in which one seeks to assign tours to vehicles in a way that satisfies operational constraints and minimizes the total cost of operations, can also be tackled by identifying the same structure in the subproblems [22], making it also amenable to analog protocols on NAQC.
@@
Figure 1: Workflow of the tested hybrid classical-quantum column generation approach for the fleet assignment problem, which uses a neutral atom quantum computer (NAQC) as a sampler of many, good, and diverse potential variables to add to the classical workflow. In the top left, a fleet assignment problem is represented as a conflict graph; nodes represent available tours, colors compatible vehicle classes, and the cost table represents the tour costs (vehicle costs not shown). After building the RMP with an initial set of columns $\Lambda^{\prime}$, one obtains its optimal dual solution, which serves to construct a series of maximum weighted independent set problems (called pricing subproblems, PSPs) that are each sent to the NAQC. There, we use our register and pulse design strategies to taylor the dynamics in order to measure a set of good and diverse independent sets for the PSPs. These are post-processed, and then the ones passing a threshold are accepted in the column set $\Lambda^{\prime}$. The process terminates when no improving column is found. An example of the primal solution after a single CG iteration is shown under ’(Re)Build RMP’.
To tackle these questions, we sought a hard CO problem on which to perform benchmarks. We chose the fleet assignment problem for its added benefit of being highly relevant to the logistics industry. We construct a QCG framework inspired from Ref. [21, 22] for the fleet assignment problem. The workflow is shown in Fig. 1, and described in detail in Sec. III.
Expanding on the analog compilation introduced in [21], we introduce a new classical algorithm for embedding a given MWIS graph instance into an spatial arrangement of atoms (a register) and a Rydberg blockade radius. Our algorithm based on simulated annealing leads to embedded graphs that have fewer extra edges and are thus more easily realized on the hardware. We also propose a new pulse design which favors quantum fluctuations towards the end of the annealing and thus increases diversity in the sampled bitstrings. We validate these insights on random graph instances. We show later on our benchmarks on synthetic instances that these changes to the analog compilation, which aim to increase the overall quality and diversity of the samples obtained, result in higher performances.
@@
The paper is organized as follow. After introducing the fleet assignment problem in Sec. II, we cover the column generation (CG) decomposition for this problem in Sec. III, as well as several classical approaches to the subproblems. In Sec. IV, we cover the fundamentals of neutral atom quantum computers, introducing our novel register embedding algorithm (Sec. IV.1) and pulse design (Sec. IV.2). Our results are presented in Sec. V, with a particular emphasis on the performance of the different protocols for the subproblems with respect to quality and diversity of the columns returned (Sec. V.4). Finally, we offer concluding remarks and final insights into the conditions under which this quantum workflow may yield a practical advantage.
@@ II Fleet Assignment
To better grasp the relationship between these objects, we introduce a conflict graph $G=(K,E)$, where each node $k\in K$ corresponds to a tour as previously defined. An edge $e_{kk^{\prime}}\in E$ is added between two nodes $k$ and $k^{\prime}$ if 1) the corresponding tours cannot be assigned to the same vehicle due to overlapping time windows ($(t^{k}_{\text{start}}<t^{k^{\prime}}_{\text{end}})\wedge(t^{k}_{\text{end}}<t^{k^{\prime}}_{\text{start}})$) or 2) incompatible vehicle classes ($V_{k}\cap V_{k^{\prime}}=\emptyset$). Other constraints could lead to more edges being added to the conflict graph. Therefore, any subset of the nodes $\tilde{K}\subset K$ such that not two nodes of $\tilde{K}$ share an edge is a feasible vehicle-to-tour assignment. This is exactly the definition of an independent set on the conflict graph $G$. We represent an independent set as the binary vector $\vec{a}=\{a_{1},...,a_{|K|}\}$, where $a_{k}=1$ denotes the presence of the route $k$ in the independent set and 0 denotes its absence. We denote $\Omega(G)$ as the set of all independent sets in $G$. For a given independent set $\vec{a}$, one also has to assign the compatible vehicle fleet $v$.
We denote a specific assignment with another binary vector $\vec{b}=\{b_{1},...,b_{|V|}\}$ where $b_{v}=1$ if $v$ is the vehicle type chosen and 0 otherwise, such that $\sum_{v=1}^{|V|}b_{j}=1$, i.e. only one vehicle class has been assigned to this set of routes. Together, the set of all vehicle-to-tour assignments, denoted by $\Lambda$, can formally be expressed as:
@@
The challenge of solving the fleet assignment problem is evident from the problem stated above. The size of the set $\Lambda$ increases exponentially in both the number of vehicle classes $|V|$ and of tours to service $|K|$. Thus, navigating such a large set of potentially feasible solutions is the bottleneck that makes brute-force approaches impractical. In practical settings, this problem is often modeled as a mixed-integer linear program (MILP) (see Ref. [22] for an example formulation), where the number of assigned vehicles per class is an unbounded positive natural number. State-of-the-art approaches to MILPs include branch-and-bound, column and row generation and bender’s decompositions [33, 34]. These techniques are sometimes further improved by greedy heuristics. The software Gurobi [35] implements most of these approaches and is among the best black-box solvers of MILPs, thus we use it to benchmark the hardness of our tested instances. In the rest of the paper, we explore the column generation framework to solve the fleet conversion problem, and describe how this leads to a series of subproblems to solve that are especially compatible with analog neutral atom quantum computers (NAQCs).
@@ III.2 Pricing Sub-Problem
Once a solution to the LP is obtained with the limited set $\Lambda^{\prime}$, one can formulate the pricing sub-problem (PSP) whose solution will lead to new columns to augment $\Lambda^{\prime}$. The key constraints for the formulation of the PSP are those from the dual formulation of Eq. (5). They can be combined to give:
@@
Note that $\mu_{k}$, $\mu_{v}^{\min}$ and $\mu_{v}^{\max}$ are the dual solutions to the LP. Thus, the goal of CG is then to find new variables $\lambda_{s^{\prime}}=(\vec{b}_{s^{\prime}},\vec{a}_{s^{\prime}})$ such that they violate the inequality shown above. The addition of these new columns in the primal problem amounts to the addition of new constraints in the dual problem, will lead to a reduction of the optimal cost of the dual problem and thus to a lower overall minimum for the fleet assignment problem. The PSP then seeks to find such variables by solving the following maximization problem:
@@
Any solution tuple $(\vec{m}^{\ast},\vec{n}^{\ast})$ with cost $\sigma^{\ast}$ where $\sigma^{\ast}>0$ then represents a new violation of the inequality of Eq. (6). It is possible, and in fact likely, that multiple such columns exist. It is then added to the set of columns $\Lambda^{\prime}\rightarrow\{\lambda_{\ast}\}\bigcup\Lambda^{\prime}$.
@@ IV Analog Quantum Algorithm for Diversified Pricing
In the presence of the drive with a maximum Rabi $\Omega_{\max}=\max_{t}\Omega(t)$, a pair of atoms will undergo a dynamical effect called the Rydberg blockade [53]. This effectively prohibits the quantum state from having any weight in the $|r_{1},r_{2}\rangle=|1,1\langle$ computational state. If two atoms are closer than the Rydberg blockade radius $R_{b}$, given by:
@@ IV.1 Embedding Strategy
By tuning $\lambda$ in Eq. (14), one can penalize more strongly missing or extra edges. In NAQCs, limits on $R_{B}$ and on the minimum atomic distance make it so that there is an upper limit to the number of edges per node that can be practically realized. Thus, for general graphs with a number of edge per node commensurate with $|V|$, it is guaranteed that the embedding will have missing edges. In this case, one can use a fast greedy postprocessing step that explores other bitstrings that are $O(1)$ Hamming distance away from the original. For MWIS problems, this postprocessing seeks to A) fix any bitstring provided to make it independent (i.e. reintroduce missing edges) and B) greedily add nodes to the proposed independent set if they can increase the total weight (see details in Appendix B). On the other hand, extra edges in $G_{H}$ (which represent independence constraints that are not present in the original graph) leads to feasible solutions being cut from our (classical or quantum) sampling procedure. We therefore use $\lambda=2$ in SA-embedder to promote missing edges over extra edges in our embedding.

### v2

# Leveraging Analog Neutral-Atoms Quantum Computers for Diversified Pricing in Hybrid Column Generation Frameworks

@@ I Introduction
Combinatorial optimization (CO) problems involve selecting the optimal solution from a finite, yet exponentially large set of possible solutions subject to constraints. They are ubiquitous in domains such as logistics, scheduling, and resource allocation, where even marginal improvements in optimization can yield substantial operational and economic benefits. Quantum algorithms have thus attracted attention for their potential to speedup certain classes of these problems [1, 2, 3, 4, 5, 6, 7], which would amplify the operational and economic gains. Nonetheless, current industrial implementations are rare.
@@
While quantum computing offers alternative algorithmic strategies based on entanglement and superposition, it operates in a large Hilbert space of size $O(2^{N})$, with $N$ the number of qubits in a quantum system. Quantum algorithms must therefore manipulate quantum information in this space, crafting interference patterns so as to concentrate probability on the answer. An example of this recipe for unstructured search is Grover’s algorithm [3], which offers a quadratic speedup over a brute force search [8], though additionnal limitations may be present [9]. Such digital algorithms for CO mostly remain out of reach in the noisy intermediate-scale quantum (NISQ) regime due to the prohibitively large number of coherent operations they require.
@@
Quantum annealing approaches, inspired by the adiabatic theorem [10, 11, 12], have been explored extensively as more NISQ-friendly protocols [13, 14, 15, 16]. They typically start by representing an optimization problem as a quadratic unconstrained binary optimization (QUBO), where constraints become slack variables and Lagrange multipliers [17, 18]. Through the annealing, one carries a known product state under a time-varying Hamiltonian such that the final Hamiltonian encodes the cost function (i.e. the QUBO). In the adiabatic limit, the final state will be the solution to the QUBO, although that limit is generally unachievable for NISQ devices with a finite coherence time. Instead, annealing is done in a finite time, and one typically encounters a very small energy gap which leads to the proliferation of defects [19, 20]. Even in a fully coherent system, these defects may lead to observed states that are not even feasible solutions [21]. The quantum approximate optimization algorithm (QAOA) [22, 23], a heuristic method designed to approximate optimal annealing paths on digital quantum computers, suffers from similar problems, although methods such as subspace embedding [24, 25] can remedy them, at the cost of deeper circuits.
@@
Classical algorithms, such as branch-and-price-and-cut [26], harness the structure of the constrained optimization problem for increased performance. They use it to refine heuristics, find cuts, generate new variables, set new bounds. On the other hand, quantum approaches such as those presented above rely mostly on unstructured search. In this dichotomy lies the need for quantum algorithms that take advantage of efficient mathematical formulations, while tackling the biggest bottlenecks with new quantum protocols. Quantum column generation (QCG) schemes [27, 28, 29, 30, 31] fill this need. They leverage the classical column generation (CG) framework, where the problem is split into an easily solvable master problem and hard subproblems with a common structure. The subproblems focus on identifying new columns (or variables) that may improve the master problem’s optimum. These subproblems are themselves hard CO problems that are solved iteratively many times by heuristics, representing a major bottleneck for performance. Initial results suggest that quantum approaches to the subproblems can lead to better overall results than other heuristics, such as genetic algorithms [28], simulated annealing [30, 27], and even exact methods [30, 27].
@@
In some particular cases, such as the graph coloring problem, the subproblems’ structure as maximum weighted independent sets (MWIS) instances makes them particularly amenable to neutral-atoms quantum computers (NAQC) [27]. This is because the analog and continuous control of these interacting many-body systems can reach a dynamical regime dominated by the Rydberg blockade, which forbids atoms closer than a certain distance from being both in their respective excited states. This is exactly the independence constraint on edges in a MWIS instance: $n_{i}+n_{j}\leq 1$ ($n_{i}$ is a projector on the Rydberg excited state). Another practically important CO problem, the fleet assignment problem, in which one seeks to assign tours to vehicles in a way that satisfies operational constraints and minimizes the total cost of operations, can also be tackled by identifying the same structure in the subproblems [28], making it also amenable to analog protocols on NAQC.
@@
Figure 1: Workflow of the tested hybrid classical-quantum column generation approach for the fleet assignment problem, which uses a neutral-atoms quantum computer (NAQC) as a sampler of many, good, and diverse potential variables to add to the classical workflow. In the top left, a fleet assignment problem is represented as a conflict graph; nodes represent available tours, colors compatible vehicle classes, and the cost table represents the tour costs (vehicle costs not shown). After building the RMP with an initial set of columns $\Lambda^{\prime}$, one obtains its optimal dual solution, which serves to construct a series of maximum weighted independent set problems (called pricing subproblems, PSPs) that are each sent to the NAQC. There, we use our register and pulse design strategies to tailor the dynamics in order to measure a set of good and diverse independent sets for the PSPs. These are post-processed, and then the ones passing a threshold are accepted in the column set $\Lambda^{\prime}$. The process terminates when no improving column is found. An example of the primal solution after a single CG iteration is shown under ’(Re)Build RMP’.
To tackle these questions, we sought a hard CO problem on which to perform benchmarks. We chose the fleet assignment problem for its added benefit of being highly relevant to the logistics industry. We construct a QCG framework inspired from Ref. [27, 28] for the fleet assignment problem. The workflow is shown in Fig. 1, and described in detail in Sec. III. Expanding on the analog compilation introduced in [27], we introduce a new classical algorithm for embedding a given MWIS graph instance into an spatial arrangement of atoms (a register) and a Rydberg blockade radius. Our algorithm based on simulated annealing leads to embedded graphs that have fewer extra edges and are thus more easily realized on the hardware. We also propose a new pulse design which favors quantum fluctuations towards the end of the annealing and thus increases diversity in the sampled bitstrings. We validate these insights on random graph instances. We show later on our benchmarks on synthetic instances that these changes to the analog compilation, which aim to increase the overall quality and diversity of the samples obtained, result in higher performances.
@@
The paper is organized as follows. After introducing the fleet assignment problem in Sec. II, we cover the column generation (CG) decomposition for this problem in Sec. III, as well as several classical approaches to the subproblems. In Sec. IV, we cover the fundamentals of neutral-atoms quantum computers, introducing our novel register embedding algorithm (Sec. IV.1) and pulse design (Sec. IV.2). Our results are presented in Sec. V, with a particular emphasis on the performance of the different protocols for the subproblems with respect to quality and diversity of the columns returned (Sec. V.4). Finally, we offer concluding remarks and final insights into the conditions under which this quantum workflow may yield a practical advantage.
@@ II Fleet Assignment
To better grasp the relationship between these objects, we introduce a conflict graph $G=(K,E)$, where each node $k\in K$ corresponds to a tour as previously defined. An edge $e_{kk^{\prime}}\in E$ is added between two nodes $k$ and $k^{\prime}$ if 1) the corresponding tours cannot be assigned to the same vehicle due to overlapping time windows ($(t^{k}_{\text{start}}<t^{k^{\prime}}_{\text{end}})\wedge(t^{k}_{\text{end}}<t^{k^{\prime}}_{\text{start}})$) or 2) incompatible vehicle classes ($V_{k}\cap V_{k^{\prime}}=\emptyset$). Other constraints could lead to more edges being added to the conflict graph. Therefore, any subset of the nodes $\tilde{K}\subset K$ such that not two nodes of $\tilde{K}$ share an edge is a feasible vehicle-to-tour assignment. This is exactly the definition of an independent set on the conflict graph $G$. We represent an independent set as the binary vector $\vec{a}=\{a_{1},...,a_{|K|}\}$, where $a_{k}=1$ denotes the presence of the route $k$ in the independent set and 0 denotes its absence. We denote $\Omega(G)$ as the set of all independent sets in $G$. For a given independent set $\vec{a}$, one also has to assign the compatible vehicle fleet $v$. We denote a specific assignment with another binary vector $\vec{b}=\{b_{1},...,b_{|V|}\}$ where $b_{v}=1$ if $v$ is the vehicle type chosen and 0 otherwise, such that $\sum_{v=1}^{|V|}b_{j}=1$, i.e. only one vehicle class has been assigned to this set of routes. Together, the set of all vehicle-to-tour assignments, denoted by $\Lambda$, can formally be expressed as:
@@
The challenge of solving the fleet assignment problem is evident from the problem stated above. The size of the set $\Lambda$ increases exponentially in both the number of vehicle classes $|V|$ and of tours to service $|K|$. Thus, navigating such a large set of potentially feasible solutions is the bottleneck that makes brute-force approaches impractical. In practical settings, this problem is often modeled as a mixed-integer linear program (MILP) (see Ref. [28] for an example formulation), where the number of assigned vehicles per class is an unbounded positive natural number. State-of-the-art approaches to MILPs include branch-and-bound, column and row generation and bender’s decompositions [40, 41]. These techniques are sometimes further improved by greedy heuristics. The software Gurobi [42] implements most of these approaches and is among the best black-box solvers of MILPs, thus we use it to benchmark the hardness of our tested instances. In the rest of the paper, we explore the column generation framework to solve the fleet conversion problem, and describe how this leads to a series of subproblems to solve that are especially compatible with analog neutral-atoms quantum computers (NAQCs).
@@ III.2 Pricing Sub-Problem
Once a solution to the LP is obtained with the limited set $\Lambda^{\prime}$, one can formulate the pricing sub-problem (PSP) whose solution will lead to new columns to augment $\Lambda^{\prime}$. The formulation of the PSP is derived from the constraints of the dual formulation of Eq. (5), which can be combined to give:
@@
where that $\mu_{k}$, $\mu_{v}^{\min}$ and $\mu_{v}^{\max}$ are the dual solutions to the LP. The goal of CG is to find new columns $\lambda_{s^{\prime}}=(\vec{b}_{s^{\prime}},\vec{a}_{s^{\prime}})$ for which the above inequality does not hold, indicating that the current solution is not yet optimal. The addition of these new columns in the primal problem amounts to the addition of new constraints in the dual problem. Therefore, from the perspective of the dual problem, the PSP searches for a column $\lambda_{s^{\prime}}$ whose coefficients $a_{k,s^{\prime}}$ and $b_{v,s^{\prime}}$ maximize the violation of the above inequality (i.e. maximize the left-hand side) under the current dual solution ($\mu_{k}$, $\mu_{v}^{\min}$, $\mu_{v}^{\max}$). These coefficients also need to respect the constraints of the primal problem (Eq. (4)). Thus, the PSP can be written as a maximization problem:
@@
where we replace coefficients $a_{k,s^{\prime}}$ and $b_{v,s^{\prime}}$ by $n_{k}$ and $m_{v}$ respectively. Note that we drop the subscript $s^{\prime}$ since the PSP focuses on finding a single candidate column. Any solution tuple $(\vec{m}^{\ast},\vec{n}^{\ast})$ with cost $\sigma^{\ast}$ where $\sigma^{\ast}>0$ then represents a new violation of the inequality of Eq. (6). It is possible, and in fact likely, that multiple such columns exist. It is then added to the set of columns $\Lambda^{\prime}\rightarrow\{\lambda_{\ast}\}\bigcup\Lambda^{\prime}$.
@@ IV Analog Quantum Algorithm for Diversified Pricing
In the presence of the drive with a maximum Rabi $\Omega_{\max}=\max_{t}\Omega(t)$, a pair of atoms will undergo a dynamical effect called the Rydberg blockade [60]. This effectively prohibits the quantum state from having any weight in the $|r_{1},r_{2}\rangle=|1,1\rangle$ computational state. If two atoms are closer than the Rydberg blockade radius $R_{b}$, given by:
@@ IV.1 Embedding Strategy
By tuning $\lambda$ in Eq. (14), one can penalize more strongly missing or extra edges. In NAQCs, limits on $R_{B}$ and on the minimum atomic distance make it so that there is an upper limit to the number of edges per node that can be practically realized. Thus, for general graphs with a number of edge per node commensurate with $|V|$, it is guaranteed that the embedding will have missing edges. In this case, one can use a fast greedy postprocessing step that explores other bitstrings that are $O(1)$ Hamming distance away from the original. For MWIS problems, this postprocessing seeks to A) fix any bitstring provided to make it independent (i.e. reintroduce missing edges) and B) greedily add nodes to the proposed independent set if they can increase the total weight (see details in Appendix B). On the other hand, extra edges in $G_{H}$ (which represent independence constraints that are not present in the original graph) leads to feasible solutions being cut from our (classical or quantum) sampling procedure. We therefore use $\lambda=2$ in SA-embedder to promote missing edges over extra edges in our embedding. This choice represents a practical balance as choosing much larger values of $\lambda$ tends to produce an embedding graph that is increasingly fragmented and sparse, making the quantum evolution on the resulting graph trivial to simulate classically.
@@
A brief comment on how embedding quality with SA-embedder may scale beyond the tested sizes is useful. Since the embedding problem is typically NP-hard for general graphs, the number of updates affecting $\mathcal{O}(1)$ atoms to reach the optimal value loss function value $C_{\lambda}$ is expected to scale exponentially with system size unless the graph has a trivial structure (e.g., a 1D chain, fully disconnected, or fully connected graphs). Moreover, the optimal value $C_{\lambda}$ is expected to grow as $N$ increases due to the increasing number of potential edges $\mathcal{O}(pN^{2})$, where $p$ is the edge probability, that can cause UDG violations. Note that the dependence on $p$ is non-trivial: while increasing $p$ generally increase the optimal $C_{\lambda}$, the extreme case $p=1$ (complete graph) has a perfect UDG representation. To improve embedding quality for larger systems, one could consider: 1) an enhanced annealing schedule with an increased number of iterations, and 2) cluster updates in which $\mathcal{O}(N)$ atoms are moved or swapped simultaneously. In this context, cluster moves appear particularly promising for improving embedding quality, since in larger graphs, we found that the nodes naturally tend form separate disconnected clusters at high $\beta$, and connecting them would require moving entire clusters close to each other.